# D. Leykekhman, B. Vexler and J. Wagner

### v1

@@ 8.2 Example 1
Here, ${\bf{y}}$ has been constructed in such a way, that $\nabla\cdot{\bf{y}}=0$ and $\|{\bf{y}}\|_{L^{2}(\Omega)}=1$. It was obtained by considering the potential $\rho(x_{1},x_{2})=\left(\sin(\pi x_{1})\sin(\pi x_{2})\right)^{4}$, and normalizing the vector field $(\partial_{x_{2}}\rho(x_{1},x_{2}),-\partial_{x_{1}}\rho(x_{1},x_{2}))^{T}$. We choose $\bar{{\bf{u}}}=\varphi(t){\bf{y}}(x_{1},x_{2})$, ${\bf{w}}={\bf{y}}$ and $\beta\equiv 1$. It is then straightforward to verify, that $G_{\bf{w}}(\bar{{\bf{u}}})\leq\beta$ for all $t\in I$ and $G_{\bf{w}}(\bar{{\bf{u}}})=\beta$ if and only if $t\in[1/4,3/4]$. We thus choose the multiplier $\bar{\mu}=10^{3}\chi_{[1/4,3/4]}(t)$, which by construction satisfies $\bar{\mu}\geq 0$ and $\langle G_{\bf{w}}(\bar{\bf{u}})-\beta,\bar{\mu}\rangle=0$. We then proceed by choosing $\bar{p}=0$ and as a consequence set $\bar{\bf{q}}=\partial_{t}\bar{\bf{u}}-\Delta\bar{\bf{u}}$. We obtain $\bar{\bf{z}}=-\bar{\bf{q}}$ and with the choice $\bar{r}=0$ can fix ${\bf{u}}_{d}=\bar{\bf{u}}+\partial_{t}\bar{\bf{z}}+\Delta\bar{\bf{z}}+\bar{\mu}{\bf{w}}$, in such a way, that the constructed $(\bar{\bf{q}},\bar{\bf{u}})$ satisfy the first order optimality condition for this desired state. Note that $\vec{\psi}$ has been chosen sufficiently smooth at the boundary, such that $\Delta\vec{\psi}|_{\partial\Omega}=0$ and thus $\bar{\bf{q}}|_{\partial\Omega}=\bar{\bf{z}}|_{\partial\Omega}=0$. The calculation of the analytic solution was verified using the SageMath software [33]. We discretize this problem with a uniform triangulation of $\Omega$ and a uniform partition of $I$. To get more insight into the observed orders of convergence, for sequences of discretization levels $\{h_{l}\}$,$\{k_{l}\}$ we report the empirical orders of convergence determined by
$(\mathrm{EOC}_{h})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l}h_{l}}\|_{L^{2}(I;L^{2}(\Omega))})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l-1}h_{l-1}}\|_{L^{2}(I;L^{2}(\Omega))})}{\log(h_{l})-\log(h_{l-1})}$
$(\mathrm{EOC}_{k})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l}h_{l}}\|_{L^{2}(I;L^{2}(\Omega))})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l-1}h_{l-1}}\|_{L^{2}(I;L^{2}(\Omega))})}{\log(k_{l})-\log(k_{l-1})}$
@@
Figure 1(a) displays convergence with respect to the spacial discretization parameter for fixed $k$. The theoretical convergence order of 1 can be observed. Figure 1(b) depicts the convergence results for the time discretization parameter for different values of $h$. Note that for this comparison, we have choosen discretization levels in time, such that the two boundary of the active set $[1/4,3/4]$ are midpoints of the discrete subintervals, in order to exclude superconvergence effects. Due to the structure of the discretization, the number of degrees of freedom, grows rather quickly, which is why very fine discretizations are expensive. Figure 1(b) shows, that if the spacial discretization parameter is chosen too large, the stagnation phase sets in rather early. If one only observes the coarse discretizations, the observed order of convergence is skewed, which led to an estimated order of convergence of 0.85 reported in [7]. As the derivation of the present example with an analytic reference solution requires the involved functions to have some precise regularity, we next propose an example with desired state that has low regularity in time. In this case a numerical reference solution is needed to measure the errors.
@@
((a)) Convergence with respect to $h$ for $k=10^{-3}$.
@@
((b)) Convergence with respect to $k$ for different values of $h$.
Figure 1: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$ for Example 1.
@@ 8.3 Example 2
and the scalar constraint is given by $\beta\equiv 1$. Moreover, we consider the regularization parameter $\alpha=10^{-4}$. As in this case, no analytical optimal solution is known, we estimate the errors using a numerical reference solution on a fine grid. To this end, we discretize the problem on 960 time intervals and 128 subdivisions of $\Omega$ in each direction. Note that due to this evaluation of the errors, we expect a faster convergence than theoretically derived, as on the finest discretization level, the error would equal to 00. In Figure 3(a) we can again observe order 1 convergence with respect to $h$ for fixed time discretization. In Figure 3(b) we observe that the convergence with respect to $k$ exhibits a rate of about $0.6$, which is much closer to the analytically derived $0.5$ than the rate observed for the smooth example 1.
((a)) Time function $\varphi(t)$.
((b)) Weight function ${\bf{w}}$.
((c)) Satisfaction of the state constraint.
@@
((a)) Convergence with respect to $h$ for $k=960^{-1}$.
@@
((b)) Convergence with respect to $k$ for $h=2^{-7}$.
Figure 3: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$ for Example 2.

### v2

@@ 8.2 Example 1
Here, ${\bf{y}}$ has been constructed in such a way, that $\nabla\cdot{\bf{y}}=0$ and $\|{\bf{y}}\|_{L^{2}(\Omega)}=1$. It was obtained by considering the potential $\rho(x_{1},x_{2})=\left(\sin(\pi x_{1})\sin(\pi x_{2})\right)^{4}$, and normalizing the vector field $(\partial_{x_{2}}\rho(x_{1},x_{2}),-\partial_{x_{1}}\rho(x_{1},x_{2}))^{T}$. We choose $\bar{{\bf{u}}}=\varphi(t){\bf{y}}(x_{1},x_{2})$, ${\bf{w}}={\bf{y}}$ and $\beta\equiv 1$. It is then straightforward to verify, that $G_{\bf{w}}(\bar{{\bf{u}}})\leq\beta$ for all $t\in I$ and $G_{\bf{w}}(\bar{{\bf{u}}})=\beta$ if and only if $t\in[1/4,3/4]$. We thus choose the multiplier $\bar{\mu}=10^{3}\chi_{[1/4,3/4]}(t)$, which by construction satisfies $\bar{\mu}\geq 0$ and $\langle G_{\bf{w}}(\bar{\bf{u}})-\beta,\bar{\mu}\rangle=0$. We then proceed by choosing $\bar{p}=0$ and as a consequence set $\bar{\bf{q}}=\partial_{t}\bar{\bf{u}}-\Delta\bar{\bf{u}}$. We obtain $\bar{\bf{z}}=-\bar{\bf{q}}$ and with the choice $\bar{r}=0$ can fix ${\bf{u}}_{d}=\bar{\bf{u}}+\partial_{t}\bar{\bf{z}}+\Delta\bar{\bf{z}}+\bar{\mu}{\bf{w}}$, in such a way, that the constructed $(\bar{\bf{q}},\bar{\bf{u}})$ satisfy the first order optimality condition for this desired state. Note that $\vec{\psi}$ has been chosen sufficiently smooth at the boundary, such that $\Delta\vec{\psi}|_{\partial\Omega}=0$ and thus $\bar{\bf{q}}|_{\partial\Omega}=\bar{\bf{z}}|_{\partial\Omega}=0$. The calculation of the analytic solution was verified using the SageMath software [37]. We discretize this problem with a uniform triangulation of $\Omega$ and a uniform partition of $I$. To get more insight into the observed orders of convergence, for a sequence of discretization levels $\{\sigma_{l}\}=\{(h_{l},k_{l})\}$, we report the empirical orders of convergence determined by
$(\mathrm{EOC}_{h})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l-1}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})}{\log(h_{l})-\log(h_{l-1})}$
$(\mathrm{EOC}_{k})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l-1}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})}{\log(k_{l})-\log(k_{l-1})}$
@@
Figure 3 displays convergence with respect to the spacial discretization parameter for fixed $k$. The theoretical convergence order of 1 can be observed. Figure 3 depicts the convergence results for the time discretization parameter for different values of $h$. Note that for this comparison, we have choosen discretization levels in time, such that the two boundary of the active set $[1/4,3/4]$ are midpoints of the discrete subintervals, in order to exclude superconvergence effects. Due to the structure of the discretization, the number of degrees of freedom, grows rather quickly, which is why very fine discretizations are expensive. Figure 3 shows, that if the spacial discretization parameter is chosen too large, the stagnation phase sets in rather early. If one only observes the coarse discretizations, the observed order of convergence is skewed, which led to an estimated order of convergence of 0.85 reported in [8]. As the derivation of the present example with an analytic reference solution requires the involved functions to have some precise regularity, we next propose an example with desired state that has low regularity in time. In this case a numerical reference solution is needed to measure the errors.
@@
Figure 1: Convergence with respect to $h$ for $k=10^{-3}$.
@@
Figure 2: Convergence with respect to $k$ for different values of $h$.
Figure 3: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ for Example 1.
@@ 8.3 Example 2
and the scalar constraint is given by $\beta\equiv 1$. Moreover, we consider the regularization parameter $\alpha=10^{-4}$. As in this case, no analytical optimal solution is known, we estimate the errors using a numerical reference solution on a fine grid. To this end, we discretize the problem on 960 time intervals and 128 subdivisions of $\Omega$ in each direction. Note that due to this evaluation of the errors, we expect a faster convergence than theoretically derived, as on the finest discretization level, the error would equal to 00. In Figure 10 we can again observe order 1 convergence with respect to $h$ for fixed time discretization. In Figure 10 we observe that the convergence with respect to $k$ exhibits a rate of about $0.6$, which is much closer to the analytically derived $0.5$ than the rate observed for the smooth example 1.
Figure 4: Time function $\varphi(t)$.
Figure 5: Weight function ${\bf{w}}$.
Figure 6: Satisfaction of the state constraint.
@@
Figure 8: Convergence with respect to $h$ for $k=960^{-1}$.
@@
Figure 9: Convergence with respect to $k$ for $h=2^{-7}$.
Figure 10: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ for Example 2.
8.4 Example 3
In the previous examples, no control constraints were present. To highlight, that the derived error estimates are indeed not influenced by the control constraints, we augment the Example of Section 8.2 by a control constraint $\bar{\bf{q}}\leq 200\cdot\mathds{1}$. The remaining choices of $\Omega,{\bf{w}},{\bf{u}}_{d},\alpha$ and $\beta$ are kept the same. Due to the presence of the control constraint, an analytic solution $\bar{\bf{q}}$ to the optimal control problem is not known, and we again compare to a fine-grid solution, computed with $k=960^{-1}$ and $h=2^{-7}$. The numerically observed orders of convergence in $k$ and $h$, as well as the active set of the control contstraint for a fixed poin in time can be observed in Figure 11. It has to be noted, that while the discrete problem has a similar structure to the one without control constraint, and can be solved by the same PDAS algorithm, the resulting saddle point systems can be much larger for large active sets. Efficient preconditioners for such problems are needed, to study the performance of this algorithm for finer discretizations.
$k=960^{-1}$
$h$ err $\mathrm{EOC}_{h}$
$2^{-2}$ 64.049 -
$2^{-3}$ 35.438 0.85
$2^{-4}$ 16.024 1.14
$2^{-5}$ 6.8962 1.21
$h=2^{-7}$
$k$ err $\mathrm{EOC}_{k}$
$5^{-1}$ 18.915 -
$10^{-1}$ 9.1295 1.05
$20^{-1}$ 4.5816 0.99
$40^{-1}$ 2.3133 0.98
$80^{-1}$ 1.1640 0.99
$160^{-1}$ 0.5827 0.99
Figure 11: Left: Orders of convergence for an example with control constraints. Right: Active sets of the discrete optimal control for $h=2^{-5}$, $k=30^{-1}$ at $t=1/3$. (grey: inactive, blue: constraint in $x_{1}$ direction active, red: constraint in $x_{2}$ direction active).